OpenStudy (anonymous):

given the original equation,determine the value(s) of a variable i n which there are... one solution, 2 com conj solutions, a neg solution. i dont have examples, but can someone tell me what i should do?

OpenStudy (anonymous):

for quadratics

OpenStudy (anonymous):

For quadratics, how can we get only one solution?? We get two solutions, that can be same or equal, is that what you mean?

OpenStudy (anonymous):

i think it's asking for discriminant

OpenStudy (anonymous):

and there's also a question like : given the discrimants, find the original equation?

OpenStudy (anonymous):

my teacher never taught this..

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

here's an example : value{s) of k for which there is only one solution 16x^2+kx+9=0

OpenStudy (anonymous):

See, in an equation of the form : \(ax^2 + bx + c =0\) Discriminant (D) is given by: \[D = b^2 - 4ac\]

OpenStudy (anonymous):

i know how to find this one, but how would i find values that make it have complex conjugates, a neg solution, or whatever else

OpenStudy (anonymous):

Can you find its Discriminant?

OpenStudy (anonymous):

yes i did that^ so the answer for ONE solution would be +- 24

OpenStudy (anonymous):

the discrimant is-552 or -600

OpenStudy (anonymous):

+-24 to make it a perfect square

OpenStudy (anonymous):

i mean 0

OpenStudy (anonymous):

for one solution

OpenStudy (anonymous):

how would i find values to make it a negative solution... for example

OpenStudy (anonymous):

See : \(D = b^2 - ac = k^2 - 4 \cdot 16 \cdot 9 = 0 = k^2 = 576\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So, if \(k= \pm 24\), then you will get one solution..

OpenStudy (anonymous):

\[D = 0 \implies One \ solution \ or \ Equal \ Solution \\ D > 0 \implies Two \ Distinct \ \mathbb{R} \ Solutions \\ D < 0 \implies Two \ Complex \ Solutions\]

OpenStudy (anonymous):

For two complex conjugate solutions, D should be less than 0..

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

like an inequality

OpenStudy (anonymous):

\[\implies k^2 - 576 < 0 \\ k^2 < 576 \\ k < \pm 24\]

OpenStudy (anonymous):

oh, that's more simpler than i thought! lol

OpenStudy (anonymous):

Choose any value of \(k\) less than 24 but greater than -24..

OpenStudy (anonymous):

but how would i find for 2 unequal real solutions?

OpenStudy (anonymous):

\(-24 < k < 24\) This will give you two complex..

OpenStudy (anonymous):

For that D > 0..

OpenStudy (anonymous):

wouldn't negative solution and complex be the same?

OpenStudy (anonymous):

So, \(k^2 > 576\), this will give you two unequal solutions..

OpenStudy (anonymous):

how come^?

OpenStudy (anonymous):

yes, negative and two complex are same, if D is negative, then two complex solutions you will get..

OpenStudy (anonymous):

\[D > 0 \\ k^2 - 576 > 0 \\ k^2 > 576\]

OpenStudy (anonymous):

By this you will get two unequal real solutions for D > 0..

OpenStudy (anonymous):

wouldn't that simplify to k>+-24?

OpenStudy (anonymous):

So, for \(k >24\) and \(k< -24\) you will get two unequal real solutions..

OpenStudy (anonymous):

2 unequal real solutions is the same as 2real irrational right?

OpenStudy (anonymous):

real irrational?? Real are the bigger set of numbers, it includes both rational and irrational numbers..

OpenStudy (anonymous):

wouldnt k>+-24 bring out to ---> k>-24 and K>24?

OpenStudy (anonymous):

why would it be k<-24 and k>24

OpenStudy (anonymous):

2 unequal real solutions or two distinct solutions we say.. But yes, if solution is like : \(2+ \sqrt{3}\) and \(2 - \sqrt{3}\), these are real irrational here..

OpenStudy (anonymous):

Choose a number lesser than -24..

OpenStudy (anonymous):

I choose : \(k=-25\).. Find D for it, is it coming positive?

OpenStudy (anonymous):

uhyes

OpenStudy (anonymous):

ah okay, i guess that's how it is..

OpenStudy (anonymous):

thanks, my whole class didn't understand the test question..

OpenStudy (anonymous):

So, \(k<-24\) will give you positive D.. Similarly, \(k>24\) will give you positive D. Squaring of \(k\) is making that to happen that way, getting?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it makes sense :D

OpenStudy (anonymous):

If you have doubt anywhere, then just take a value for \(k\) then check for D. If it is 0, then one solution. If D < 0, then two complex if D > 0, then two unequal real..

OpenStudy (anonymous):

All good?

OpenStudy (anonymous):

i also have like 2 more questions D:

OpenStudy (anonymous):

that i want to clear up

OpenStudy (anonymous):

Sure, go ahead.. :)

OpenStudy (anonymous):

so if you were to be given a discriminant, is it possible to find the original equations?

OpenStudy (anonymous):

I will arrange help for you as I don't have much time to stay here more, don't worry you will get help here always..

OpenStudy (anonymous):

cause it's possible if you were given the roots

OpenStudy (anonymous):

With only given D, it is difficult to say..

OpenStudy (anonymous):

ah okay. thanks !!!

OpenStudy (anonymous):

i think that's all :o

OpenStudy (anonymous):

Do you have any example with you?

OpenStudy (anonymous):

@apoorvk or @iGreen can we have your eyes here? Not physically though..!! :P

OpenStudy (anonymous):

well i wasn't sure, because my friend said that there were questions like : he was given a discrimant and had to find the original equation... but i wasn't sure if he was given extra info...

OpenStudy (anonymous):

but he probably over exaggerated

OpenStudy (anonymous):

You need it now or you can wait? If nothing comes here to help you then I can but tomorrow at the same time or sometime earlier than this, but tomorrow..

OpenStudy (anonymous):

Is that okay>

OpenStudy (anonymous):

nah i have until tomorrow ::)

OpenStudy (anonymous):

but thanks so much. if you could figure it out. tell me!

OpenStudy (anonymous):

Sure.. :)

OpenStudy (anonymous):

Check this links to practice questions: http://2012books.lardbucket.org/books/beginning-algebra/s12-04-guidelines-for-solving-quadrat.html

OpenStudy (anonymous):

See, I have broken my head on your question, but according to me, if only D is given, you cannot figure out its Original Equation..

OpenStudy (anonymous):

Because for you can get same D for two or more equations..